OpenStudy (yb1996):

At what angle (in degrees) does first-order diffraction from layers of atoms 161.0 pm apart occur, using x rays with a wavelength of 147 pm?

OpenStudy (jadedry):

You have to use Bragg's diffraction equation: \[n*\lambda = 2d*\sin \theta\] where n = the order of the diffraction. In this case, it equals 1. Lamba = 147 pm or 1.47 * 10^10 m D = 161 pm or 1.61 *10^10m Sin theta is the variable you have to solve for. \[1 * 1.47*10^{10} = 1.61*10^{10} *2 * \sin \theta\] \[0.45652 = \sin \theta\] Angle of diffraction = 27.16 degrees

OpenStudy (yb1996):

Thank You!!!